# Paper Quality Assessment based on Individual Wisdom Metrics from Open Peer Review

### v1

@@ Abstract
This study proposes a data-driven framework for enhancing the accuracy and efficiency of scientific peer review through an open, bottom-up process that estimates reviewer quality. Traditional closed peer review systems, while essential for quality control, are often slow, costly, and subject to biases that can impede scientific progress. Here, we introduce a method that evaluates individual reviewer reliability by quantifying agreement with community consensus scores and applying Bayesian weighting to refine paper quality assessments. We analyze open peer review data from two major scientific conferences, and demonstrate that reviewer-specific quality scores significantly improve the reliability of paper quality estimation. Perhaps surprisingly, we find that reviewer quality scores are unrelated to authorship quality. Our model incorporates incentive structures to recognize high-quality reviewers and encourage broader coverage of submitted papers, thereby mitigating the common “rich-get-richer” pitfall of social media. These findings suggest that open peer review, with mechanisms for estimating and incentivizing reviewer quality, offers a scalable and equitable alternative for scientific publishing, with potential to enhance the speed, fairness, and transparency of the peer review process.
@@ 1 Introduction
Scientific communication, the primary focus of the present work, suffers from related problems. Partially for historical reasons (Birukou et al.,, 2011) and partly to head off the kinds of misinformation problems found in open media, scientific review is a closed, slow, top-down, expensive process with incentives for high-profile publishing that might also distort science (Noorden,, 2013; Young et al.,, 2008; Buranyi,, 2017; da Silva and Dobránszki,, 2014). A scientific report is submitted to one journal, whose editors select whether to desk-reject it or send it out for review to a small number of hand-picked reviewers. The resulting reviews emanate from a small number of reviewers and are thus highly stochastic. If the paper is rejected, the process repeats until the paper is accepted at some journal. This friction-filled process slows scientific progress in a way that can impede junior scientists’ careers. The exponentially growing volume of scientific output makes these challenges more daunting. The concept of ”publish then review”, with bottom-up review, offers a potential antidote (Eisen et al.,, 2020; Ginsparg,, 1997; Eisen,, 2016; Kravitz and Baker,, 2011; Kriegeskorte et al.,, 2012; Nosek and Bar-Anan,, 2012; Stern and O’Shea,, 2019; da Silva,, 2013; Walther and van den Bosch,, 2012). Submitted papers are immediately considered as published, and review is an ongoing process that begins after publication. This process requires making reviewing open, equitable, and bottom-up, so that any scientist can review any paper, by self-selection (LeCun,, 2013). However, how to make such an open system reliable, and with the right incentive mechanisms, is an open problem (Sandewall,, 2012; Birukou et al.,, 2011; Pöschl,, 2012; Goldberg et al.,, 2024).
In this paper, we take a data-driven approach to the question of how to extract a reliable signal from noisy and flawed data provided by the scientific community. We first quantify the level of agreement between reviewers in peer review data, and show that it is surprisingly small. Next, we propose a method for assessing paper quality based on an estimate of reviewer quality for each individual. Interestingly, reviewership quality is independent of authorship quality. We show that this metric can support an open, democratic peer-review solution that tackles the following challenges: It allows reviewers to self-organize by picking the works that they want to review. It provides recognition and thus incentives to reviewers for consistently writing high-quality reviews. In addition, it allows an effective extraction of the community belief about the quality of papers from scarce and noisy reviews.
@@ 2.1 High variability: A core challenge for collective paper review
We consider data from two conferences that employed a transparent review process and used explicit scoring for determining paper acceptance decisions: ICLR 2023 and the Conference on Cognitive Computational Neuroscience (CCN) 2023. In the CCN2023 peer review process, each of the 589 reviewers assigned ”Impact” and ”Clarity” numerical scores to the 527 submitted abstracts. All reviewers were authors of submitted abstracts. Each abstract received an average of 9.1 reviews. In ICLR2023, each of the 3798 submitted papers received anywhere from 2 to 9 reviews by reviewers selected through a separate track, with an average of 3.8 reviews per submission. Each review consisted of multiple dimensions of scores, with the overall rating and confidence scores (used in analyses in this paper).
@@
In both datasets, the correlation in scores given by pairs of reviewers to the same paper was surprisingly low (Figure 1). The correlation was very weak in the community-based review process of CCN ($r=0.161\pm 0.014$; Figure 1A). Though somewhat higher for ICLR2023 ($r=0.361\pm 0.014$; Figure 1C), the correlations were still small despite the fact that the submissions being assessed were full-length standalone papers. Common possible causes of low correlation, such as variations in what range of scores each reviewer actually utilizes, are sometimes addressed by score normalization techniques (e.g., z-scoring, ranking, mean removal, or distribution inversion). Applying any of these methods only marginally increased the correlation ($r<0.19$ for CCN2023; Figure 1B). Although the correlation increased with the self-reported confidence scores in ICLR2023 data (from $r=0.24$ for confidence=2 up to $r=0.49$ for confidence=5; Figure 1D), only $10.0\%$ of reviewers reported having this high confidence. Moreover, the confidence of ratings was inversely correlated with the ratings scores (Supplementary Figure S1), implying that on average, better papers receive less confident reviews. In sum, our analysis reveals a striking lack of agreement among reviewers when assessing the same paper.
@@
2.2 Bayesian paper quality estimation using reviewer quality estimatates
@@
Figure 2: In a simple model, weighting the reviews using a reviewer quality metric results in a better estimation of paper quality. (A) The simple model: different reviewers all assign scores to the same paper according to a distribution centered at the ground truth quality (hidden), but with different standard deviations, corresponding to different reviewer accuracies. (B) Bayes-optimal weighting of review scores leads to a tighter estimation of the ground truth paper quality scores (reviewer accuracies either taken from ground truth or estimated from five previous reviews). Shaded regions denote the s.d.
@@
We first consider a very simple model (Figure 2; a slightly richer model is presented below). Assume that each submission (paper) $j$ has a hidden ground truth quality $q_{j}$. Reviewers are assumed to give scores to that paper drawn from a distribution centered at $q_{j}$, with a user-specific standard deviation $\sigma_{i}$, where $i$ is the index of the reviewer (Figure 2A). We can think of the inverse standard deviation of a reviewer as a ground-truth reviewer quality measure. The goal is to obtain good estimates of $q_{j}$ without knowledge of the paper ground truth values. This simple model leaves out possible systematic reviewer biases (see Shah et al., (2017); Helmer et al., (2017); Ross, (2017); Stelmakh et al., (2019); Goldberg et al., (2024); Stelmakh et al., (2023) for discussion of biases in the review process), assumes there are no bots or bad actors (we add bots below), and doesn’t constrain the scores to lie in a bounded interval, however it suffices to demonstrate the core idea. Traditionally, the scores of all reviewers are averaged to obtain a final estimated quality score of the paper, on which decisions are made. This calculation produces an estimator of the paper quality score with mean squared deviation (MSD) equal to
[…]
@@
If reviewer quality is known, the optimal weighting technique, a Bayesian approach, assigns weights to reviewers’ scores that are proportional to the inverse square of the estimated standard deviation of that reviewer (for derivation of these results, see Appendix 1): $w_{i}\propto 1/\sigma_{i}^{2}$. This approach leads to a paper quality estimator with
[…]
@@
This Bayesian paper quality estimator based on known reviewer quality much more tightly follows the ground truth quality of the paper (Figure 2B, black) than the standard mean (Figure 2B, gray).
In the real setting, reviewers’ standard deviations are not known a priori, requiring that they are estimated from the data. The Bayesian MSD using an empirically estimated reviewer quality still significantly outperforms the standard mean (Figure 2B, black dashed line), suggesting that using estimates of individual reviewer quality can improve estimates of paper quality.
2.3 Assessing the Bayesian approach in CCN 2023 review data
@@
Figure 3: The reviewer quality metric, unrelated to authorship quality, improved community score estimation of CCN2023 submissions. (A) In the CCN2023 review process, reviewer quality (measured as MSD of user’s review scores from community average score across all reviewed papers) and author quality (measured as average score the user got for their paper) were unrelated. To preserve reviewer privacy in this visualization, some outlier points were removed from the plot and reviewers were binned together in small bins. (B) Users with higher estimated reviewer quality (Bayes weight), when estimating paper quality, consistently had smaller deviations from community score on withheld papers. Shaded regions denote the s.e.m. (C) Weighting by the inferred Bayes-optimal coefficient resulted in a significantly reduced mean squared deviation from the community score on withheld papers. By contrast, weighting by the authorship quality (”author score”) did not produce benefits. Error bars denote the s.e.m.
The above result motivates applying a Bayesian weighting of paper reviews by estimated reviewer quality in a real world review process. To assess outcomes, we again turn to the CCN2023 process (Figure 3). In the experiments that follow, we used the CCN2023 ”Impact” scores.
Unfortunately, in such real world datasets we lack ground truth knowledge of paper quality to assess the robustness of different measures. To construct a proxy, we assume that the full community average score (CAS) for a paper is the ground truth. The CAS represents the score that would be assigned if many scientists in the community were to review the paper. With an estimate for individual reviewer quality, we can assess the efficacy of different metrics on subsampled sets of reviews for each paper.
CCN2023 data contained unique reviewer IDs for each reviewer, across papers, thus we were able to compute reviewer standard deviations relative to the CAS and generate an estimate of reviewer quality. Unlike in our simple model, scientists in CCN2023 play two roles: they are reviewers and authors. A quality value can be assigned to each role. Given the sparsity of the data (each individual writes relatively few reviews and each paper is reviewed by relatively few reviewers), a natural question is whether quality as an author is related to quality as a reviewer and thus whether it might be possible to arrive at a better per-individual reviewership score by combining both.
@@
We considered the mean squared deviation of an individual’s scores from the CAS each paper reviewed by that individual (i.e., their quality as a reviewer), and compared it with the CAS they received for their own submissions (i.e., their quality as an author). This resulted in the surprising discovery that authorship quality is unrelated to reviewership quality (Figure 3A; $r=-0.023,p=0.57,N=539$ reviewers). The finding challenges the assumption that high-scoring authors are necessarily good at predicting the community judgment of a paper’s quality. Instead, there were reviewers who received very high scores for their own submissions yet were remarkably poor at predicting the community judgment of papers they reviewed (Figure 3A, inset, upper right). Conversely, some reviewers who received much lower scores for their own submissions were able to estimate the CAS quality of reviewed papers much more accurately (Figure 3A, inset, lower right). In fact, the group of reviewers that was able to estimate the CAS with the highest accuracy was the group with intermediate authorship scores (Figure 3, Supplementary Figure S2A-B), and this group was also had the highest level of agreement between reviewers reviewing the same paper (Figure 3, Supplementary Figure S3; $r=0.352$ for the middle percentiles, N=795 pairs). This discrepancy between the authorship score and the reviewer quality might be attributable to various factors, including a lack of rewards and recognition for putting effort into conducting a thorough review.
In sum, the reviewer quality metric is independent of the author quality metric, and author quality likely cannot be used as a way to augment reviewer quality scores.
@@
Improving paper assessments through empirical reviewer quality estimates
@@
Individuals with a higher reviewer quality score consistently demonstrated a better ability to predict the CAS score. This is a cross-validated result, shown for withheld papers – those not included in the estimation of the users’ reviewer quality scores, thus avoiding information leakage (Figure 3B).
All reviewers can be combined to obtain a better prediction of the CAS than simple averaging or averaging with traditional normalizers, even when working with a limited number of reviews (Figure 3C; from $MSD=0.0554\pm 0.001$ for the simple mean estimator to $MSD=0.0464\pm 0.001$ for the Bayes weighting procedure). On the contrary, using the author quality score instead of the reviewer quality score does not lead to a significant improvement ($MSD=0.0553\pm 0.001$).
This improved estimation technique has the potential to enhance the overall quality and fairness of the peer review process, particularly when the number of reviews per paper is limited. We were only able to test this hypothesis in the CCN 2023 dataset, because the ICLR open process does not provide a consistent reviewer identifier across the user’s reviews and submissions, which made it impossible to aggregate data on the performance of the reviewer across more than just one review.
2.4 How things could be: an open framework for peer review
Figure 4: Benefits of reviewer quality metrics in a richer ecosystem. (A) The conception of an open platform, where content is published immediately, and scored by reviewers that are self-selected from the user pool, and assessed on the review quality by the same community. (B) Binning reviewers by rating scores allows the estimation of the accuracy of any group in assessing the community belief about the quality of any given submission. (C) Bayes weighting of review scores based on the estimated reviewer accuracy metric from (B) leads to a better estimate of the ground truth paper quality scores (left; solid black line) than simple averaging of all review scores (left; gray line). An alternative heuristic of simply thresholding out all users below a certain rating percentile (here, below 80%, red line) results in a similar correlation to Bayes weighting but lower coverage of papers on the platform (right). (D) Dsitributions of estimated reviewer quality scores for bots and real users in the simulated platform. In high-noise regime with most of the users being bots (up to 80%), reliability of the reviewer quality estimation breaks down, suggesting that alternative ways of excluding bots might be more effective.

### v2

# Paper scoring through individual reviewer quality estimates for Open Peer Review

@@ Abstract
Traditional closed peer review systems, which have played a central role in scientific publishing, are often slow, costly, non-transparent, stochastic, and possibly subject to biases – factors that can impede scientific progress and undermine public trust. Here, we propose and examine the efficacy and accuracy of an alternative form of scientific peer review: through an open, bottom-up process. First, using data from two major scientific conferences (CCN2023 and ICLR2023), we highlight how high variability of review scores and low correlation across reviewers presents a challenge for collective review. We quantify reviewer agreement with community consensus scores and use this as a reviewer quality estimator, showing that surprisingly, reviewer quality scores are not correlated with authorship quality. Instead, we reveal an inverted U-shape relationship, where authors with intermediate paper scores are the best reviewers. We assess empirical Bayesian methods to estimate paper quality based on different assessments of individual reviewer reliability. We show how under a one-shot review-then-score scenario, both in our models and on real peer review data, a Bayesian measure significantly improves paper quality assessments relative to simple averaging. We then consider an ongoing model of publishing, reviewing, and scoring, with reviewers scoring not only papers but also other reviewers. We show that user-generated reviewer ratings can yield robust and high-quality paper scoring even when unreliable (but unbiased) reviewers dominate. Finally, we outline incentive structures to recognize high-quality reviewers and encourage broader reviewing coverage of submitted papers. These findings suggest that a self-selecting open peer review process is potentially scalable, reliable, and equitable with the possibility of enhancing the speed, fairness, and transparency of the peer review process.
@@ 1 Introduction
Scientific communication, the primary focus of the present work, suffers from related problems. Partially for historical reasons (Birukou et al.,, 2011) and partly to head off the kinds of misinformation problems found in open media, scientific review is a closed, slow, top-down, expensive process with incentives for high-profile publishing that might also distort science (Noorden,, 2013; Young et al.,, 2008; Buranyi,, 2017; da Silva and Dobránszki,, 2014). A scientific report is submitted to one journal, whose editors select whether to desk-reject it or send it out for review to a small number of hand-picked reviewers. The resulting reviews emanate from a small number of reviewers and are thus highly stochastic. If the paper is rejected, the process repeats until the paper is eventually accepted at some journal. This friction-filled process slows scientific progress in a way that can impede junior scientists’ careers. The exponentially growing volume of scientific output makes these challenges more daunting. The concept of ”publish then review”, with bottom-up review, offers a potential antidote (Eisen et al.,, 2020; Ginsparg,, 1997; Eisen,, 2016; Kravitz and Baker,, 2011; Kriegeskorte et al.,, 2012; Nosek and Bar-Anan,, 2012; Stern and O’Shea,, 2019; da Silva,, 2013; Walther and van den Bosch,, 2012). Submitted papers are immediately considered as published, and review is an ongoing process that begins after publication. This process requires making reviewing open, equitable, and bottom-up, so that any scientist can review any paper by self-selection (LeCun,, 2013). However, how to make such an open system reliable, and with the right incentive mechanisms, is an open problem (Sandewall,, 2012; Birukou et al.,, 2011; Pöschl,, 2012; Goldberg et al.,, 2024).
In this paper, we take a data-driven approach to the question of how to extract a reliable signal from noisy and flawed data provided by the scientific community. We first quantify the level of agreement between reviewers in peer review data, and show that it is surprisingly small. Next, we propose a method for assessing paper quality based on an estimate of reviewer quality for each individual. Interestingly, reviewership quality is independent of authorship quality. We show that this metric can support an open, democratic peer-review solution that tackles the following challenges: it allows reviewers to self-organize by picking the works that they want to review; it provides recognition and thus incentives to reviewers for consistently writing high-quality reviews; in addition, it allows an effective extraction of the community belief about the quality of papers from scarce and noisy reviews.
@@ 2.1 High variability: A core challenge for collective paper review
We consider data from two conferences that employed a transparent review process and used explicit scoring for determining paper acceptance decisions: the International Conference on Learning Representations (ICLR) 2023 and the Conference on Cognitive Computational Neuroscience (CCN) 2023. In the CCN2023 peer review process, each of the 589 reviewers assigned ”Impact” and ”Clarity” numerical scores to the 527 submitted abstracts. All reviewers were authors of submitted abstracts. Each abstract received an average of 9.1 reviews. In ICLR2023, each of the 3798 submitted papers received anywhere from 2 to 9 reviews by reviewers selected through a separate track, with an average of 3.8 reviews per submission. Each review consisted of multiple dimensions of scores, with the overall rating and confidence scores used in analyses in this paper.
@@
In both datasets, the correlation in scores given by pairs of reviewers to the same paper was surprisingly low (Figure 1). The correlation was very weak in the community-based review process of CCN ($r=0.161\pm 0.014$; Figure 1A). Though somewhat higher for ICLR2023 ($r=0.361\pm 0.014$; Figure 1C), the correlations were still small despite the fact that the submissions being assessed were full-length standalone papers. Common possible causes of low correlation, such as variations in what range of scores each reviewer actually utilizes, are sometimes addressed by score normalization techniques (e.g., z-scoring, ranking, mean removal, or distribution inversion). Applying any of these methods only marginally increased the correlation ($r<0.19$ for CCN2023; Figure 1B). Although the correlation increased with the self-reported confidence scores in ICLR2023 data (from $r=0.24$ for confidence=2 up to $r=0.49$ for confidence=5; Figure 1D), only $10.0\%$ of reviewers reported having this high confidence. Moreover, the confidence of ratings was inversely correlated with the paper quality rating scores (Supplementary Figure S1), implying that on average, better papers receive less confident reviews. In sum, our analysis reveals a striking lack of agreement among reviewers when assessing the same paper.
@@
2.2 Bayesian paper quality estimation using empirical reviewer quality estimates
@@
Figure 2: In a simple model, weighting reviews using an empirical reviewer quality metric results in better estimation of paper quality. (A) The simple model: different reviewers all assign scores to the same paper according to a distribution centered at the ground truth quality (hidden), but with different standard deviations, corresponding to different reviewer accuracies. (B) Bayesian weighting of review scores based on reviewer quality leads to a tighter estimation of the ground truth paper quality scores. Curves shown for ground truth reviewer quality (black solid line), empirically estimated reviewer quality based on five previous reviews (black dashed line), and simple mean score (solid gray line). Shaded regions denote the s.d.
Methods: Estimating paper quality using reviews.
@@
To demonstrate how a potential remedy to this challenge can look like, we first consider a very simple model (Figure 2; a slightly richer model is presented below). Assume that each submission (paper) $j$ has a hidden ground truth quality $q_{j}$. Reviewers are assumed to give scores to that paper drawn from a distribution centered at $q_{j}$, with a user-specific standard deviation $\sigma_{i}$, where $i$ is the index of the reviewer (Figure 2A). We can think of the inverse standard deviation of a reviewer as a ground-truth reviewer quality measure. The goal is to obtain good estimates of $q_{j}$ without knowledge of the paper ground truth values. This simple model leaves out possible systematic reviewer biases (see Shah et al., (2017); Helmer et al., (2017); Ross, (2017); Stelmakh et al., (2019); Goldberg et al., (2024); Stelmakh et al., (2023) for discussion of biases in the review process), assumes there are no bots or bad actors (we add bots below), and doesn’t constrain the scores to lie in a bounded interval, however it suffices to demonstrate the core idea. Traditionally, the scores of all reviewers are averaged to obtain a final estimated quality score of the paper, the score on which decisions are made. This calculation produces an estimator of the paper quality score with mean squared deviation (MSD) equal to
$\text{MSD(simple mean)}=\frac{\sum\sigma_{i}^{2}}{n^{2}}.$ (1)
@@
However, if reviewer quality is known, the optimal weighting technique (a Bayesian approach) is to assign weights to reviewers’ scores that are proportional to the inverse square of the estimated standard deviation of that reviewer (for derivation of these results, see Appendix 1): $w_{i}\propto 1/\sigma_{i}^{2}$. This approach leads to a paper quality estimator with a strictly lower MSD (unless all $\sigma_{i}$ are the same):
$\text{MSD(Bayesian)}=\frac{1}{\sum 1/\sigma_{i}^{2}}.$ (2)
@@
This Bayesian paper quality estimator based on known reviewer quality follows the ground truth quality of the paper much more tightly (Figure 2B, solid black) than the standard mean (Figure 2B, gray).
In real settings, the reviewers’ quality is of course not known a priori, requiring that this parameter be estimated from the data. To simulate this situation, we compute an empirically estimated reviewer quality based on five previous reviews by each reviewer, by using the mean squared deviation of those reviews from the community average paper score. In this process, we assume that the community score was collected from enough other reviewers to estimate the ground truth qualities of those papers with high accuracy and no bias. When we apply Bayesian paper quality estimation with the empirically determined reviewer quality estimates, we find that the strategy still significantly outperforms the standard mean (Figure 2B, black dashed line). Our simple model setting, therefore, suggests that estimating individual reviewer quality and applying those quality estimates in a Bayesian way to papers can can improve estimates of paper quality. Next, we test these ideas in on real-world conference data.
2.3 Applying the Bayesian approach to CCN 2023 review data
@@
Figure 3: Reviewer quality, unrelated to author quality, improved paper quality estimation in CCN2023 submissions. (A) In the CCN2023 review process, reviewer quality (measured as MSD of user’s review scores from community average score across all reviewed papers) and author quality (measured as the average score the user got for their paper) were unrelated. To preserve reviewer privacy in this visualization, some outlier points were removed from the plot and reviewers were binned together in small bins. (B) Using estimated author quality as a stand-in for reviewer quality as the Bayes weight did not improve the estimate of paper quality relative to baseline (cross-validated results, shown on withheld papers). Error bars denote the s.e.m. (C) Users with higher estimated reviewer quality (Bayes weight estimated on the training data) consistently had smaller deviations from community score on withheld papers when estimating reviewer quality. Shaded regions denote the s.e.m. (D) Using the estimated reviewer quality as Bayes weights led to consistently smaller deviations from the community score on withheld papers. Error bars denote the s.e.m.
To assess outcomes in real-world settings, we turn to the data from CCN2023 (Figure 3). CCN2023 data contained unique reviewer IDs for each reviewer, across papers, and a unique author ID for each author across their submitted papers. Thus, we were able to consider different methods to compute reviewer quality scores and determine the efficacy of using those scores in paper quality estimation.
Methods: Proxy for ground-truth paper quality.
The first challenge in using real world datasets is that we lack ground-truth knowledge of paper quality, hampering our ability to compare different estimators of paper quality. Under the assumption that if a large fraction of the community were to assess a paper, the average estimate would be the ground truth measure of quality, we construct a proxy measure for ground-truth: the community average score (CAS) for each paper. There are various important discussions about the true quality of a scientific paper, such as the durability of its findings over time; however, even large-scale expert review cannot estimate durability because it involves complex future prediction. Thus our focus, thoughout this paper, is on the question of how one can arrive at a reasonable estimate of the CAS from a severely limited number of noisy reviews.
Below, we compute the CAS for each paper as the average of the “impact” score assigned by all reviewers of that paper. We then assess the efficacy of a paper quality estimation technique by applying it to subsampled sets of reviews for the paper, and comparing the estimate to the CAS.
@@
Scientists in CCN2023 play two roles, unlike in our simple model: they are both reviewers and authors. A quality value can be assigned to each role. A natural question is whether the two properties interact: for instance, is being a good author predictive of being a good reviewer? If they are related, then given the sparsity of the data (each individual writes relatively few reviews and each paper is reviewed by relatively few reviewers), combining estimates for quality in both roles might enable a more robust and better reviewer quality estimate.
@@
To answer this question, we considered the mean squared deviation of an individual’s scores from the CAS for each paper reviewed by that individual (i.e., their quality as a reviewer), and compared it with the CAS they received for their own submissions (i.e., their quality as an author). This resulted in the surprising discovery that authorship quality is unrelated to reviewership quality (Figure 3A; r=0.023,p=0.57,N=539r=-0.023,p=0.57,N=539 reviewers). This finding challenges the assumption that high-scoring authors are necessarily good at predicting the community judgment of a paper’s quality. Instead, there were reviewers who received very high scores for their own submissions yet were remarkably poor at predicting the community judgment of papers they reviewed (Figure 3A, inset, upper right). Conversely, some reviewers who received much lower scores for their own submissions were able to estimate the CAS quality of reviewed papers much more accurately (Figure 3A, inset, lower right). In fact, the group of reviewers that was able to estimate the CAS with the highest accuracy was the group with intermediate authorship scores (Figure 3, Supplementary Figure S2A-B), and this group was also had the highest level of agreement between reviewers reviewing the same paper (Figure 3, Supplementary Figure S3; $r=0.352$ for the middle percentiles, N=795 pairs). This discrepancy between the authorship score and the reviewer quality might be attributable to various factors, including a lack of rewards and recognition for putting effort into conducting a thorough review. Thus, at least with reviews conducted as they were in CCN2023, reviewer quality appears to be independent of author quality, and author quality likely cannot be used in a direct way as a measure of reviewer quality.
We test this hypothesis as follows. The baseline method for paper quality estimation as a function of number of reviewers is simple averaging of subsets of reviewer scores. We compare this baseline against a method which weighs the reviewer’s scores using the authorship quality of each reviewer, using empirically estimated author quality as a measure of reviewer quality. Authorship quality is estimated by the average of the CAS on all papers of that author. Using this empirical authorship quality measure as weights led to no measurable improvement in the MSD around the CAS of the resulting paper quality estimates relative to simple mean estimators, Figure 3B ($MSD=0.0553\pm 0.001$ versus $MSD=0.0554\pm 0.001$ for the mean estimator). Thus, we conclude that (estimated) author quality is not directly proportional to (estimated) reviewer quality and does not, applied in a simple way, result in better paper quality estimation.
In the results of this section, the cross-validation was done in a leave-one-out fashion: that is, to estimate the deviation of a reviewer’s score of a paper from the CAS, that reviewer was temporarily removed from the dataset, the CAS was calculated without it, and then compared to that reviewer’s score to obtain the deviation.
@@
Improved CCN 2023 assessments through empirical Bayesian reviewer quality weighting
We next consider whether and how direct reviewer quality estimates can improve paper quality estimation, according to the procedures of our simple model above (Figure 2). To do so, for every paper in the dataset, we estimate its reviewers’ Bayes weights by leaving this paper out of the dataset, then calculating the mean squared deviation of the reviewers’ scores from CAS on all the other papers, and using the formula $1/\sigma_{\text{estimated}}^{2}$ to obtain the Bayes weights for each reviewer according to Eq. 2. We were able to perform these tests in the CCN 2023 dataset but not the ICLR data because ICLR’s process does not provide a consistent reviewer identifier across reviewed papers.
@@
To obtain a cross-validated result that avoids information leakage between the data used to obtain reviewer quality scores and the data on which we evaluate paper quality, we only consider performance on withheld papers – those not included in the estimation of the users’ reviewer quality scores. We find that indeed, individuals with a higher estimated reviewer quality score consistently demonstrated a better cross-validated ability to predict the CAS score (Figure 3C).
Given this result, we next generate a cross-validated prediction of the CAS by combining paper scores from reviewers based on their reviewer quality scores. We find that this procedure yields a better prediction of the CAS than simple averaging of paper scores (Figure 3D): performance is $MSD=0.046\pm 0.001$, compared to a higher $MSD=0.055\pm 0.001$ with the simple mean estimator ($p<0.001$).
In sum, this estimation process – which first estimates reviewer reliability based on a subset of their reviews and the CAS, then performs Bayesian weighting of paper scores – has the potential to enhance the overall quality and fairness of peer reviews when the number of reviews per paper is limited (in this dataset, the mean number of reviews per paper is 9). Next, we consider an alternative method for estimating reviewer quality, based on direct peer evaluation of reviews, and compare the two methods.
2.4 Open peer review with user-generated ratings of both papers and reviews
Building on the findings above, we hypothesize that a new open framework for scientific peer review may be feasible and scalable, involving bottom-up self-selection of papers to review by users, and users scoring others’ reviews (Figure 4).